OpenStudy (lena772):

The function below shows the number of students of a school who enrolled for salsa classes. Let f(x) represent the total number of students who enrolled for the classes after x years. f(x) = 10(1.18)^x The average rate of change in the number of students who enrolled for salsa classes from the second to the fifth year is ________students per year. Round your answer to the nearest whole number.

OpenStudy (the_fizicx99):

f(5) = 10(1.18)10 f(5) = 10 * 5.9 f(5) = 59 --- f(2) = 10(1.18)2 f(2) = 10* 2.36 f(2) = 23.6 f(b) - f(a)/b-a 59 - 23.6/5-2 35.4/3 = ?

OpenStudy (lena772):

u so dumb it's ^x

OpenStudy (the_fizicx99):

Dumbpellet... put the damn ^ -.-

OpenStudy (lena772):

I did u tard

OpenStudy (anonymous):

In the second year, no of students who enrolled class is f(2) = 10 * 1.18^2 = 13.924 ~ 14 (students) Similarly, in fifth year, f(5) = 10 * 1.18^5 = 22.87 ~ 23 (students) The average rate of change from 2nd to 5th year is \[\frac{f(5)-f(2)}{5-2} = 3\] Average rate of change is 3 students per year

OpenStudy (lena772):

ty @linh412986